L^{A}TEX News

Issue 1, June 1994

Welcome to L^{A}TEX News

An issue of L^{A}TEX News will accompany every future release of L^{A}TEX. It will tell you about important events, such as major bug fixes, newly available packages, or any other L^{A}TEX news.

L^{A}TEX 2ε—the new L^{A}TEX release

The most important news is the release of L^{A}TEX 2ε, the new version of the L^{A}TEX software. This version has better support for fonts, graphics and colour, and will be actively maintained by the L^{A}TEX3 project team. Upgrades will be issued every six months, in June and December.

Why a new L^{A}TEX?

Over the years many extensions have been developed for L^{A}TEX. This is, of course, a sure sign of its continuing popularity but it has had one unfortunate result: incompatible L^{A}TEX formats came into use at different sites. Thus, to process documents from various places, a site maintainer was forced to keep L^{A}TEX (with and without NFSS), SliTEX, AMS-L^{A}TEX, and so on. In addition, when looking at a source file it was not always clear for which format the document was written.

To put an end to this unsatisfactory situation a new release of L^{A}TEX was produced. It brings all such extensions back under a single format and thus prevents the proliferation of mutually incompatible dialects of L^{A}TEX 2.09. The new release was available for several months as a test version, and the final release of 1 June officially replaces the old version.

Processing documents with L^{A}TEX 2ε

Documents written for L^{A}TEX 2.09 will still be read by L^{A}TEX 2ε. Any such document is run in L^{A}TEX 2.09 compatibility mode.

Unfortunately, compatibility mode comes with a price: it can run up to 50% slower than L^{A}TEX 2.09 did. If you want to run your document in the faster native mode, you should try replacing the line:

\documentstyle[options,packages]{class}

with:

   \documentclass[options]{class}
   \usepackage{latexsym,packages}

Unfortunately, this will not always work, because some L^{A}TEX 2.09 packages will only work in L^{A}TEX 2ε compatibility mode. You should find out if there is a L^{A}TEX 2ε version of the package available.

L^{A}TEX 2ε native mode also gives access to the new features of L^{A}TEX 2ε, described in L^{A}TEX 2ε for authors.

New packages

L^{A}TEX 2ε has much better support for graphics, colour, fonts, and multi-lingual typesetting. The following software should be available from the distributor who brought you L^{A}TEX 2ε:

babel, for typesetting in many languages.

color, for colour support. graphics, for including images. mfnfss, for using bitmap fonts. psnfss, for using Type 1 fonts. tools, other packages by the L^{A}TEX3 team.

The packages come with full documentation, and are also described in L^{A}TEX: A Document Processing System or The L^{A}TEX Companion.

Further information

More information about L^{A}TEX 2ε is to be found in:

L^{A}TEX: A Document Preparation System, Leslie Lamport, Addison Wesley, 2nd ed, 1994. The L^{A}TEX Companion, Goossens, Mittelbach and

Samarin, Addison Wesley, 1994.

The L^{A}TEX distribution comes with documentation on the new features of L^{A}TEX:

L^{A}TEX 2ε for authors, describes the new features of L^{A}TEX documents, in the file usrguide.tex. L^{A}TEX 2ε for class and package writers, describes the

new features of L^{A}TEX classes and packages, in the

                                                      file clsguide.tex.

L^{A}TEX 2ε font selection, describes the new features of L^{A}TEX fonts for class and package writers, in the file

                                                      fntguide.tex.

For more information on TEX and L^{A}TEX, get in touch with your local TEX Users Group, or the international TEX Users Group, P. O. Box 869, Santa Barbara, CA 93102-0869, USA, Fax: +1 805 963 8358, EMail: tug@tug.org.

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 1994, all rights reserved.